# DPDETR: Decoupled Position Detection Transformer for Infrared-Visible Object Detection

### v1

@@ Abstract
Infrared-visible object detection aims to achieve robust object detection by leveraging the complementary information of infrared and visible image pairs. However, the commonly existing modality misalignment problem presents two challenges: fusing misalignment complementary features is difficult, and current methods cannot accurately locate objects in both modalities under misalignment conditions. In this paper, we propose a Decoupled Position Detection Transformer (DPDETR) to address these problems. Specifically, we explicitly formulate the object category, visible modality position, and infrared modality position to enable the network to learn the intrinsic relationships and output accurate positions of objects in both modalities. To fuse misaligned object features accurately, we propose a Decoupled Position Multispectral Cross-attention module that adaptively samples and aggregates multispectral complementary features with the constraint of infrared and visible reference positions. Additionally, we design a query-decoupled Multispectral Decoder structure to address the optimization gap among the three kinds of object information in our task and propose a Decoupled Position Contrastive DeNosing Training strategy to enhance the DPDETR’s ability to learn decoupled positions. Experiments on DroneVehicle and KAIST datasets demonstrate significant improvements compared to other state-of-the-art methods. The code will be released at https://github.com/gjj45/DPDETR.
Index Terms:
Infrared-visible object detection, DETR, Decoupled learning, Feature alignment, Denosing training
@@ I Introduction
Object detection is a fundamental task in computer vision, which has been used in various practical applications, such as video surveillance, autonomous driving, and aerial object detection. With the development of deep learning, the object detection technique has made great progress [1, 2, 3]. Since these methods are mainly designed for visible images, they are still challenged by poor imaging conditions, such as low illumination, smoke, fog, and so on. Thus, infrared images are introduced into the object detection task [4, 5, 6]. Unlike visible imaging, infrared imaging cannot capture detailed information but is unaffected by illumination, smoke, and fog occlusion conditions. To achieve robust full-time object detection, infrared-visible object detection by integrating infrared and visible image complementary information has attracted extensive attention in recent years [7, 8, 9, 10, 11, 12].
@@
Figure 1: Misalignment problem in infrared-visible object detection and the detection comparison between the common method and our method. (a) Examples of modality misalignment in infrared and visible image pairs. The yellow and red boxes represent the position of the same object in the infrared and visible modalities, respectively. The yellow dashed line represents the object’s position in the infrared modality. (b) An example of w/o DP (without Decoupled Position) method where the detection results make it difficult to distinguish the same object in two modalities due to misalignment. Our method can accurately locate objects in both modalities and identify the same object with the same ID.
@@
However, the modality misalignment problem in infrared-visible object detection is a major challenge. Most feature fusion methods usually assume that infrared-visible image pairs are well-aligned. However, as shown in Fig. 1 (a), even through manual registration, precise registration is difficult because infrared-visible images often exhibit significant visual differences and are not always captured at the same timestamp [13]. As a result, the imaging objects in two modalities for the same one are usually misaligned. This could disrupt the consistency of fused feature representation of current methods, affecting network performance. The modality misalignment problem is more prominent in aerial object detection because objects are often labeled with tightly oriented bounding boxes, requiring more precise alignment of object features in the detection task.
@@
When infrared and visible image pairs are misaligned, the performance of most methods usually degrades and it is hard to obtain accurate object detection results in both modalities. Even though objects can be correctly located in one modality (reference modality), this still leads to obvious misalignment detection errors in the other modality and even shows confusing duplicate detection results. As shown in Fig. 1 (b), misalignment detection errors and confusing duplicate detection seriously hinder the recognition of the same object in two modalities, especially when objects are dense and precise localization is crucial. This phenomenon also indicates that these methods potentially select the modality that best matches the annotated bounding box as the main reference modality, which could lead to unfair treatment of both modalities during feature fusion, resulting in a bias towards the reference modality. Therefore, it is critical to accurately locate the position of the same object in both infrared and visible modalities.
@@
Recently, some approaches have tended to address the misalignment problem in infrared-visible object detection. Zhang et al. [14], Zhou et al. [15] and Yuan et al. [16, 11, 12] solve the modality misalignment problem by predicting the proposals’ offset between two modalities in the RoI head or by predicting the feature points’ offset before feature fusion. Although these methods address the alignment of object features to some extent, they cannot guarantee the perfect alignment of two modality features since these methods used a main reference modality proposal to predict the object proposal offset in the other modality directly, without fully utilizing both infrared and visible image features. Furthermore, these methods usually output the positions of objects in the reference modality but cannot simultaneously output the accurate positions and correspondences of the objects in both modalities.
@@
In this paper, we propose a novel Decoupled Position Detection Transformer method called DPDETR to achieve instance-level feature alignment and accurate position output of the same object in both modalities. Specifically, we decouple the object position in object detection into infrared and visible positions and represent an object with category, infrared position, and visible position information. We explicitly optimize these three aspects for each object by making full use of complementary features. To achieve aligned complementary feature fusion, we propose Decoupled Position Multispectral Deformable cross-attention, which performs adaptive sampling and aggregation at the decoupled reference positions of the object in infrared and visible features. Since optimizing these three types of information simultaneously is more complicated and there is a modality gap between them, we design a query decoupled structure to achieve decoupled cross-attention for fine-optimizing each type of information. Furthermore, we design Decoupled Position Contrastive Denosing Training to increase the diversity of misalignment situations in the two modalities, helping the network learn and accelerate this decoupled position optimization paradigm.
@@
  • We propose a novel method called DPDETR with Decoupled Position Multispectral Deformable cross-attention to solve modality misalignment problems and output accurate positions of each object in both modalities. To the best of our knowledge, this is the first work to decouple object positions in infrared and visible modality and simultaneously optimize them using multimodal features.
  • We design a query decoupled structure to achieve decoupled cross-attention for fine-optimizing each type of information and propose Decoupled Position Contrastive Denosing Training to assist network training. These two approaches further enhance DPDETR’s ability to decouple learning.
@@ II-B Alignment Learning in Infrared-visible Object Detection
Modality misalignment is a key problem in infrared-visible object detection. Recently, some work has been dedicated to solving this problem. Zhang et al. [14, 7] first addressed the alignment problem by predicting the shift offset of the reference proposal in another modality and fusing aligned proposal features. [16, 11] further considered the scale and angle offsets of the reference proposal to realize more accurate alignment feature fusion in aerial object detection. [12] calculates the attention value between feature points in the reference modality and another modality to achieve the fusion of misaligned object features. However, these methods simply predict the offset of another modality based on the reference modality, without fully utilizing infrared and visible features to learn the intrinsic relationship between the same object in two modalities. In contrast, our method explicitly formulates the position of each object in both infrared and visible modalities, fully utilizes the multispectral features to learn this intrinsic relationship, and outputs the accurate position of the object in both modalities.
Figure 2: The overall architecture of DPDETR. The framework consists of feature extraction and encoding modules, a Paired IoU-aware Competitive Query Selection module, a Decoupled Multispectral Decoder, and the Decoupled Position Contrastive Denosing Training strategy. The final output of DPDETR is the category of each object and its accurate position in both visible and infrared images.
@@ II-C End to End Object Detectors
In recent years, Carion et al. [34] first proposed the end-to-end object detector based on transformer called DEtection TRansformer (DETR). It views object detection as a set prediction problem and uses binary matching to directly predict one-to-one object sets during training. However, DETR has the problem of slow training convergence and many DETR variants have been proposed to address this issue. Deformable DETR [35] accelerated training convergence by predicting 2D reference points and designing a Deformable cross-attention module to sparsely sample features around reference points. Conditional DETR [36] decoupled the content and position information and proposed conditional cross-attention to accelerate training convergence. Efficient DETR [37] built a more efficient pipeline by combining dense prediction and sparse prediction. DAB-DETR [38] introduced 4D reference points to optimize the anchor boxes layer by layer. DN-DETR [39] accelerated the training process and label-matching effect by introducing query denoising training group. DINO [40] integrated the above works to build a powerful DETR detection pipeline. Considering the computational efficiency of DETR, RT-DETR [41] achieves real-time object detection by designing an Efficient Hybrid Encoder.
Some works have also applied DETR to oriented object detection. O2DETR [42] is the first work to apply DETR to oriented object detection, and AO2-DETR [43] introduced an oriented proposal generation mechanism and oriented proposal refinement module to adopt oriented object detection. ARS-DETR [44] proposed an angle-embedded Rotaed Deformable Attention module to incorporate the angle information for extracting the aligned features in oriented object detection.
Recently, a DETR-based infrared-visible object detection network called DAMSDet [45] was introduced, which fuses misaligned object features through adaptive sparse sampling. However, this method lacks explicit constraints and could fail when the misalignment situation exceeds the reference position. In contrast, our DPDETR achieves more precise feature alignment and is not affected by the degree of misalignment.
@@ III Proposed Method
The overall architecture of the proposed DPDETR is shown in Fig. 2. A pair of matched infrared and visible images are used as input. The features of each image are first extracted by the modality-specific backbone network (e.g., ResNet50 [46]), and then two modality-specific Efficient Encoders encode these features, respectively. Subsequently, the encoded features are flattened, concatenated, and input into a paired IoU-aware competitive query selection module. This module selects salient modality features that contain more accurate information on classification, visible modality position, and infrared modality position as the initial matching object queries. Next, we duplicate these matching object queries into three copies, which serve as the initial class query, initial visible position query, and initial infrared position query. These copies are then input into the Decoupled Multispectral Decoder for decoupled learning and feature alignment fusion to obtain refined decoupled matching queries. Finally, these refined matching queries map to each object’s classification, visible modality position, and infrared modality position.
Additionally, during the training stage, we implement the Decoupled Position Denosing Training strategy to enhance the network’s decoupled learning ability and to increase the diversity of misalignment situations. More details will be explained in the following subsections.
@@ III-A Paired IoU-aware Competitive Query Selection
where $L_{\text{boxvis }}$,$L_{\text{boxir }}$, and $L_{\text{cls}}$ denote visible modality bounding boxes loss, infrared modality bounding boxes loss, and classification loss, respectively. Here, $\hat{y}$ and $y$ denote prediction and ground truth, with $\hat{y}=\left\{\hat{c},\hat{b}_{v},\hat{b}_{i}\right\}$ and $y=\left\{c,b_{v},b_{i}\right\}$. In this context, $c$, $b_{v}$, and $b_{i}$ represent categories, visible modality bounding boxes, and infrared modality bounding boxes, respectively. We introduce the visible and infrared IoU score into the loss function of the classification branch to realize the consistency constraint on the classification, visible modality localization, and infrared modality localization of selected queries.
@@ III-B Decoupled Multispectral Transformer Decoder
After the Paired IoU-aware competitive query selection, we obtain a set of initialization object queries that effectively represent the object category, visible modality position, and infrared modality position. However, there is a gap in the network optimizes these three types of information. For example, category information tends to focus more on the central features of the object, while position information focuses more on the edge features [47]. Additionally, in our infrared and visible modality scenarios, category information tends to incorporate features from both modalities, whereas modality-specific position information should focus more on features specific to each modality. This gap leads to limitations in the performance of our detectors when using the vanilla decoder structure, where a single set of queries simultaneously optimizes all three types of information and shares cross-attention. Hence, we design a decoupled decoder structure with decoupled queries and decoupled cross-attention branches for our optimization tasks to eliminate the gap between these three types of information.
@@
Figure 4: Demonstration of adding noise in terms of shift, scale, and angle to a pair of ground truth boxes. Red, light green, and dark green boxes represent the infrared GT box, positive infrared denoising box, and negative infrared denoising box, respectively. Blue, light purple, and dark purple boxes represent the visible GT box, positive visible denoising box, and negative visible denoising box, respectively.
@@
Figure 5: The structure of DPCDN group and the useage of positive and negative denosing queries.
@@
III-C Decoupled Position Contrastive DeNosing Training
In our network, simultaneously optimizing the three types of object information is more complex and largely depends on the accuracy of query-to-ground-truth matching. Therefore, inspired by DN-DETR [39], we design the Decoupled Position Contrastive DeNosing (DPCDN) training strategy. This approach bypasses the matching process and directly generates denoising queries by adding noise to the ground truth category, paired bounding box positions, sizes, and angles in both modalities.
@@ III-D Matching Cost and Loss Function
where $\mathcal{L}_{\mathrm{match}}(y_{i},\hat{y}_{\sigma(i)})$ is the matching cost between ground truth $y_{i}$ and a prediction with index $\sigma(i)$. This process is computed with the Hungarian algorithm. Since we explicitly optimize the object category, visible modality position and infrared modality position, we define $y_{i}=(c_{i},b_{vis-i},b_{ir-i})$ where $c_{i}$ is the object class label, $b_{vis-i}$ and $b_{ir-i}$ are the 5D vectors that represent visible and infrared ground truth oriented box, respectively. Thus we define $\mathcal{L}_{\mathrm{match}}(y_{i},\hat{y}_{\sigma(i)})$ as $-\mathds{1}_{\{c_{i}\neq\emptyset\}}\hat{p}_{\sigma(i)}(c_{i})+\mathds{1}_{\{c_{i}\neq\emptyset\}}\mathcal{L}_{\mathrm{box}}\big{(}b_{\nu is-i},\hat{b}_{\nu is-\sigma(i)}\big{)}+\mathds{1}_{\{c_{i}\neq\emptyset\}}\mathcal{L}_{\mathrm{box}}(b_{ir-i},\hat{b}_{ir-\sigma(i)})$ to ensure that the matched queries can accurately locate the objects in both the visible and infrared modalities. Where $L_{box}$ is composed of L1 cost and PIoU [49] cost for oriented box matching cost calculation.
TABLE I: Detection results ($m$AP, in %) on DroneVehicle Dataset. The best results are highlighted in bold.
Model Car Truck Bus Van Freight-car $m$AP Modality
Faster R-CNN (OBB) [1] 79.69 41.99 76.94 37.68 33.99 54.06 RGB
RetinaNet (OBB) [48] 78.45 34.39 69.75 28.82 24.14 47.11
S2ANet [50] 79.86 50.02 82.77 37.52 36.21 57.28
Oriented R-CNN [51] 80.26 55.39 86.84 46.92 42.12 62.30
ROI Transformer [52] 61.55 55.05 85.48 44.84 42.26 61.55
Oriented RepPoints [53] 84.40 55.00 85.80 46.60 39.50 62.30
Faster R-CNN (OBB) [1] 89.68 40.95 86.32 41.21 43.10 60.27 IR
RetinaNet (OBB) [48] 88.81 35.43 76.45 32.12 39.47 54.45
S2ANet [50] 89.71 51.03 88.97 44.03 50.27 64.80
Oriented R-CNN [51] 89.63 53.92 89.15 40.95 53.86 65.50
ROI Transformer [52] 89.65 50.98 88.86 44.47 53.42 65.47
Oriented RepPoints [53] 89.90 55.60 89.10 48.10 57.60 68.00
Halfway Fusion (OBB) [19] 89.85 60.34 88.97 46.28 55.51 68.19 RGB+IR
CIAN (OBB) [54] 89.98 62.47 88.90 49.59 60.22 70.23
AR-CNN (OBB) [7] 90.08 64.82 89.38 51.51 62.12 71.58
TSFADet [16] 89.88 67.87 89.81 53.99 63.74 73.06
C2Former [12] 90.20 68.30 89.80 58.50 64.40 74.20
GAGTDet [11] 90.82 69.65 90.46 55.62 66.28 74.57
DPDETR (Ours) 90.30 78.20 90.10 64.90 75.70 79.81
@@
where $L_{cls}$ is the Paired IoU-aware classification loss, $L_{box-vis}$ and $L_{box-ir}$ are composed of L1 loss and PIoU loss for visible and infrared oriented box regression, $L_{dn}$ is Decoupled Position Contrastive Denosing Training loss, and $L_{Aux}$ is the auxiliary loss computed at the output of each decoder layer.
@@
Model DPCDN QDS $m$AP Modality
1 Our Base Detector - - 70.40 RGB

### v2

@@ Abstract
Infrared-visible object detection aims to achieve robust object detection by leveraging the complementary information of infrared and visible image pairs. However, the commonly existing modality misalignment problem presents two challenges: fusing misalignment complementary features is difficult, and current methods cannot reliably locate objects in both modalities under misalignment conditions. In this paper, we propose a Decoupled Position Detection Transformer (DPDETR) to address these issues. Specifically, we explicitly define the object category, visible modality position, and infrared modality position to enable the network to learn the intrinsic relationships and output reliably positions of objects in both modalities. To fuse misaligned object features reliably, we propose a Decoupled Position Multispectral Cross-attention module that adaptively samples and aggregates multispectral complementary features with the constraint of infrared and visible reference positions. Additionally, we design a query-decoupled Multispectral Decoder structure to address the the conflict in feature focus among the three kinds of object information in our task and propose a Decoupled Position Contrastive DeNoising Training strategy to enhance the DPDETR’s ability to learn decoupled positions. Experiments on DroneVehicle and KAIST datasets demonstrate significant improvements compared to other state-of-the-art methods. The code will be released at https://github.com/gjj45/DPDETR.
@@ I Introduction
Object detection is a fundamental task in computer vision, which has been used in various practical applications, e.g., video surveillance, autonomous driving, aerial object detection, etc. With the development of deep learning, the object detection technique has made great progress [1, 2, 3]. Since existing methods are mainly designed for visible images, they are still challenged by poor imaging conditions, such as low illumination, smoke, fog, and so on. Thus, infrared images are introduced into the object detection task [4, 5, 6]. Unlike visible imaging, infrared imaging cannot capture detailed information but is unaffected by illumination, smoke, and fog occlusion conditions. To achieve robust full-time object detection, infrared-visible object detection by integrating infrared and visible image complementary information has attracted extensive attention in recent years [7, 8, 9, 10, 11, 12, 13, 14].
@@
Figure 1: Misalignment problem in infrared-visible object detection and the detection comparison between the common method and our method. (a) Examples of modality misalignment in infrared and visible image pairs. The yellow and red boxes represent the position of the same object in the infrared and visible modalities, respectively. The yellow dashed line represents the object’s position in the infrared modality. (b) An example of w/o DP (without Decoupled Position) method where the detection results make it difficult to distinguish the same object in two modalities due to misalignment. Our method can reliably locate objects in both modalities and identify the same object with the same ID.
@@
However, most of infrared-visible object detection methods usually assume that infrared-visible image pairs are well-aligned , which usually does not hold on in practice. As shown in Fig. 1 (a), even through manual registration, precise registration is difficult because infrared-visible images often exhibit significant visual differences and are not always captured at the same timestamp [15]. As a result, the imaging objects in two modalities for the same one are usually misaligned. This could disrupt the consistency of fused feature representation of current methods, futher decreasing the detection performance.
@@
When infrared and visible image pairs are misaligned, the most methods usually could not to obtain reliable object detection results in both modalities. Even though objects can be correctly located in one modality (reference modality), this still leads to obvious misalignment detection errors in the other modality and even shows confusing duplicate detection results. As shown in Fig. 1 (b), misalignment detection errors and confusing duplicate detection seriously hinder the recognition of the same object in two modalities, especially when objects are dense and the bounding boxes are tightly constrained. This phenomenon also indicates that these methods potentially select the modality that best matches the annotated bounding box as the main reference modality, which could lead to unfair treatment of both modalities during feature fusion, resulting in a bias towards the reference modality. Therefore, it is critical to accurately locate the position of the same object in both infrared and visible modalities.
@@
Recently, a few approaches have tended to address the misalignment problem in infrared-visible object detection. Zhang et al. [16], Zhou et al. [17] and Yuan et al. [18, 13, 14] solved the modality misalignment problem by predicting the proposals’ offset between two modalities in the RoI head or by predicting the feature points’ offset before feature fusion. Although these methods address the alignment of object features to some extent, they are unable to align misaligned features in complex scenarios since these methods use a main reference modality proposal to predict the object proposal offset in the other modality directly, without fully utilizing both infrared and visible image features. Furthermore, these methods usually output the positions of objects in the reference modality but cannot simultaneously output the reliably positions and one-to-one correspondences of the objects in both modalities.
@@
In this paper, we propose a novel Decoupled Position Detection Transformer method called DPDETR to achieve instance-level feature alignment and reliably position output of the same object in both modalities. Specifically, we decouple the object position in object detection into infrared and visible positions and represent an object with category, infrared position, and visible position information. We explicitly optimize these three aspects for each object by making full use of complementary features. To achieve aligned complementary feature fusion, we propose Decoupled Position Multispectral Deformable cross-attention, which performs adaptive sampling and aggregation at the decoupled reference positions of the object in infrared and visible features. Since optimizing these three types of information simultaneously is more complicated and they have conflicting optimization requirements, we design a query decoupled structure to achieve decoupled cross-attention for fine-optimizing each type of information. Furthermore, we design Decoupled Position Contrastive DeNoising Training to increase the diversity of misalignment situations in the two modalities, helping the network learn and accelerate this decoupled position optimization paradigm.
@@
  • We propose a novel method called DPDETR with Decoupled Position Multispectral Deformable cross-attention to solve modality misalignment problems and output reliable positions of each object in both modalities. To the best of our knowledge, this is the first work to decouple object positions in infrared and visible modality and simultaneously optimize them using multimodal features.
  • We design a query decoupled structure to achieve decoupled cross-attention for fine-optimizing each type of information and propose Decoupled Position Contrastive DeNoising Training to assist network training. These two approaches further enhance DPDETR’s ability to decouple learning.
@@ II-B Alignment Learning in Infrared-visible Object Detection
Modality misalignment is a key problem in infrared-visible object detection. Recently, some work has been dedicated to solving this problem. Zhang et al. [16, 8] first addressed the alignment problem by predicting the shift offset of the reference proposal in another modality and fusing aligned proposal features. [18, 13] further considered the scale and angle offsets of the reference proposal to realize more accurate alignment feature fusion in aerial object detection. [14] calculates the attention value between feature points in the reference modality and another modality to achieve the fusion of misaligned object features. However, these methods simply predict the offset of another modality based on the reference modality, without fully utilizing infrared and visible features to learn the intrinsic relationship between the same object in two modalities. In contrast, our method explicitly defines the position of each object in both infrared and visible modalities, fully utilizes the multispectral features to learn this intrinsic relationship, and outputs the reliable position of the object in both modalities.
Figure 2: The overall architecture of DPDETR. The framework consists of feature extraction and encoding modules, a Paired IoU-aware Competitive Query Selection module, a Decoupled Multispectral Decoder, and the Decoupled Position Contrastive DeNoising Training strategy. The final output of DPDETR is the category of each object and its reliable position in both visible and infrared images.
@@ II-C End to End Object Detectors
In recent years, Carion et al. [36] first proposed the end-to-end object detector based on transformer called DEtection TRansformer (DETR). It views object detection as a set prediction problem and uses binary matching to directly predict one-to-one object sets during training. However, DETR has the issue of slow training convergence and many DETR variants have been proposed to address this issue. Deformable DETR [37] accelerated training convergence by predicting 2D reference points and designing a Deformable cross-attention module to sparsely sample features around reference points. Conditional DETR [38] decoupled the content and position information and proposed conditional cross-attention to accelerate training convergence. Efficient DETR [39] built a more efficient pipeline by combining dense prediction and sparse prediction. DAB-DETR [40] introduced 4D reference points to optimize the anchor boxes layer by layer. DN-DETR [41] accelerated the training process and label-matching effect by introducing query denoising training group. DINO [42] integrated the above works to build a powerful DETR detection pipeline. Considering the computational efficiency of DETR, RT-DETR [43] achieves real-time object detection by designing an Efficient Hybrid Encoder.
Some works have also applied DETR to oriented object detection. O2DETR [44] was the first to apply DETR in this context, while AO2-DETR [45] introduced an oriented proposal refinement module for better adaptation to oriented object detection. ARS-DETR [46] further proposed an angle-embedded Rotaed Deformable Attention module, which incorporates angle information for extracting the rotated object features in oriented object detection.
Recently, a DETR-based infrared-visible object detection network called DAMSDet [47] was introduced, which addresses the complexity and variability of the complementary characteristics between infrared and visible modalities. In contrast, our DPDETR further addresses misalignment issues to achieve more reliable feature alignment.
@@ III Proposed Method
The overall architecture of the proposed DPDETR is shown in Fig. 2. A pair of matched infrared and visible images are used as input. The features of each image are first extracted by the modality-specific backbone network (e.g., ResNet50 [48]), and then two modality-specific Efficient Encoders encode these features, respectively. Subsequently, the encoded features are flattened, concatenated, and input into a paired IoU-aware competitive query selection module. This module selects salient modality features that contain more accurate information on classification, visible modality position, and infrared modality position as the initial matching object queries. Next, we duplicate these matching object queries into three copies, which serve as the initial class query, initial visible position query, and initial infrared position query. These copies are then input into the Decoupled Multispectral Decoder for decoupled learning and feature alignment fusion to obtain refined decoupled matching queries. Finally, these refined matching queries are used to predict each object’s classification, visible modality position, and infrared modality position.
Additionally, during the training stage, we implement the Decoupled Position DeNoising Training strategy to enhance the network’s decoupled learning ability and to increase the diversity of misalignment situations. In the following subsections, we provide a detailed introduction to each module.
@@ III-A Paired IoU-aware Competitive Query Selection
where $L_{\text{boxvis }}$$L_{\text{boxir }}$, and $L_{\text{cls}}$ denote visible modality bounding boxes loss, infrared modality bounding boxes loss, and classification loss, respectively. Here, $\hat{y}$ and $y$ denote prediction and ground truth, with $\hat{y}=\left\{\hat{c},\hat{b}_{v},\hat{b}_{i}\right\}$ and $y=\left\{c,b_{v},b_{i}\right\}$. In this context, $c$, $b_{v}$, and $b_{i}$ represent categories, visible modality bounding boxes, and infrared modality bounding boxes, respectively. We introduce the visible and infrared IoU score into the loss function of the classification branch to realize the consistency constraint on the classification, visible modality localization, and infrared modality localization of selected queries.
@@ III-B Decoupled Multispectral Transformer Decoder
After the Paired IoU-aware competitive query selection, we obtain a set of initialization object queries that effectively represent the object category, visible modality position, and infrared modality position. However, there is a conflict in feature focus among these three types of information. For example, category information tends to focus more on the central features of the object, while position information focuses more on the edge features [49]. Additionally, in our infrared and visible modality scenarios, category information tends to incorporate features from both modalities, whereas modality-specific position information should focus more on features specific to each modality. This gap leads to limitations in the performance of our detectors when using the vanilla decoder structure, where a single set of queries simultaneously optimizes all three types of information and shares cross-attention. Hence, we design a decoupled decoder structure with decoupled queries and decoupled cross-attention branches for our optimization tasks to eliminate the gap between these three types of information.
Figure 4: Demonstration of adding noise in terms of shift, scale, and angle to a pair of ground truth boxes. Red, light yellow, and dark yellow boxes represent the infrared GT box, positive infrared denoising box, and negative infrared denoising box, respectively. Blue, light orange, and dark orange boxes represent the visible GT box, positive visible denoising box, and negative visible denoising box, respectively.
@@
Figure 5: The structure of DPCDN group and the useage of positive and negative denoising queries.
@@
III-C Decoupled Position Contrastive DeNoising Training
In our network, simultaneously optimizing the three types of object information is more complex and largely depends on the accuracy of query-to-ground-truth matching. Therefore, inspired by DN-DETR [41], we design the Decoupled Position Contrastive DeNoising (DPCDN) training strategy. This approach bypasses the matching process and directly generates denoising queries by adding noise to the ground truth category, paired bounding box positions, sizes, and angles in both modalities.
@@ III-D Matching Cost and Loss Function
where $\mathfrak{S}_{N}$ represents the set of all possible permutations, $\mathcal{L}_{\mathrm{match}}(y_{i},\hat{y}_{\sigma(i)})$ is the matching cost between ground truth $y_{i}$ and a prediction with index $\sigma(i)$. This process is computed with the Hungarian algorithm. Since we explicitly optimize the object category, visible modality position and infrared modality position, we define $y_{i}=(c_{i},b_{vis-i},b_{ir-i})$ where $c_{i}$ is the object class label, $b_{vis-i}$ and $b_{ir-i}$ are the 5D vectors that represent visible and infrared ground truth oriented box, respectively. Thus we define $\mathcal{L}_{\mathrm{match}}(y_{i},\hat{y}_{\sigma(i)})$ as $\mathds{1}_{\{c_{i}\neq\emptyset\}}\hat{p}_{\sigma(i)}(c_{i})+\mathds{1}_{\{c_{i}\neq\emptyset\}}\mathcal{L}_{\mathrm{box}}\big(b_{\nu is-i},\hat{b}_{\nu is-\sigma(i)}\big)+\mathds{1}_{\{c_{i}\neq\emptyset\}}\mathcal{L}_{\mathrm{box}}(b_{ir-i},\hat{b}_{ir-\sigma(i)})$ to ensure that the matched queries can reliably locate the objects in both the visible and infrared modalities. Where $L_{box}$ is composed of L1 cost and PIoU [51] cost for oriented box matching cost calculation.
TABLE I: Detection results (mmAP, in %) on DroneVehicle Dataset. The best results are highlighted in bold.
Model Car Truck Bus Van Freight-car mmAP Speed(fps) Modality
Faster R-CNN (OBB) [1] 79.69 41.99 76.94 37.68 33.99 54.06 VIS
RetinaNet (OBB) [50] 78.45 34.39 69.75 28.82 24.14 47.11 -
S2ANet [52] 79.86 50.02 82.77 37.52 36.21 57.28 -
Oriented R-CNN [53] 80.26 55.39 86.84 46.92 42.12 62.30 -
ROI Transformer [54] 61.55 55.05 85.48 44.84 42.26 61.55 -
Oriented RepPoints [55] 84.40 55.00 85.80 46.60 39.50 62.30 -
Faster R-CNN (OBB) [1] 89.68 40.95 86.32 41.21 43.10 60.27 - IR
RetinaNet (OBB) [50] 88.81 35.43 76.45 32.12 39.47 54.45 -
S2ANet [52] 89.71 51.03 88.97 44.03 50.27 64.80 -
Oriented R-CNN [53] 89.63 53.92 89.15 40.95 53.86 65.50 -
ROI Transformer [54] 89.65 50.98 88.86 44.47 53.42 65.47 -
Oriented RepPoints [55] 89.90 55.60 89.10 48.10 57.60 68.00 -
Halfway Fusion (OBB) [21] 89.85 60.34 88.97 46.28 55.51 68.19 20.4 VIS+IR
CIAN (OBB) [56] 89.98 62.47 88.90 49.59 60.22 70.23 21.7
AR-CNN (OBB) [8] 90.08 64.82 89.38 51.51 62.12 71.58 18.2
TSFADet [18] 89.88 67.87 89.81 53.99 63.74 73.06 18.6
C2Former [14] 90.20 68.30 89.80 58.50 64.40 74.20 -
GAGTDet [13] 90.82 69.65 90.46 55.62 66.28 74.57 17.8
DAMSDet (OBB) [47] 90.30 74.90 89.70 63.20 70.70 77.75 11.4
DPDETR (Ours) 90.30 78.20 90.10 64.90 75.70 79.81 7.2
@@
where $L_{cls}$ is the Paired IoU-aware classification loss, $L_{box-vis}$ and $L_{box-ir}$ are composed of L1 loss and PIoU loss for visible and infrared oriented box regression, $L_{dn}$ is Decoupled Position Contrastive DeNoising Training loss, and $L_{Aux}$ is the auxiliary loss computed at the output of each decoder layer.
@@
Model DPCDN QDS mmAP Modality
1 Our Base Detector - - 70.40 VIS